# Mining higher-order triadic interactions

### v3

@@ Abstract
Complex systems often involve higher-order interactions which require us to go beyond their description in terms of pairwise networks. Triadic interactions are a fundamental type of higher-order interaction that occurs when one node regulates the interaction between two other nodes. Triadic interactions are found in a large variety of biological systems, from neuron-glia interactions to gene-regulation and ecosystems. However, triadic interactions have so far been mostly neglected. In this article, we propose a theoretical model that demonstrates that triadic interactions can modulate the mutual information between the dynamical state of two linked nodes. Leveraging this result, we propose the Triadic Interaction Mining (TRIM) algorithm to mine triadic interactions from node metadata, and we apply this framework to gene expression data, finding new candidates for triadic interactions relevant for Acute Myeloid Leukemia. Our work reveals important aspects of higher-order triadic interactions that are often ignored, yet can transform our understanding of complex systems and be applied to a large variety of systems ranging from biology to the climate.
@@ I Introduction
Higher-order networks [1, 2, 3, 4, 5] are key to capturing many-body interactions present in complex systems. Inferring higher-order interactions [6, 7, 8, 9, 10, 11] from real, pairwise network datasets is recognised as one of the central challenges in the study of higher-order networks [2, 12], with wide applicability across different scientific domains from biology and brain research [13, 14, 15] to finance [16, 17]. When from the knowledge of a pairwise network the problem is inferring the many-body interactions, mining higher-order interactions typically involves generative models and Bayesian approaches based on network structural properties [6, 18, 7, 9, 11]. Note, however that when the inference is performed on the basis on the knowledge of the nodes dynamical state [10, 8], inferring higher-order interactions also requires dynamical considerations.
@@
Triadic interactions [19] are a fundamental type of signed higher-order interaction that are gaining increasing attention from the statistical mechanics community  [19, 20, 21, 22, 23, 24], since they are not reducible to hyperedges or simplices. A triadic interaction occurs when one or more nodes regulate the interaction between two other nodes. The regulator nodes may either enhance or inhibit the interaction between the other two nodes. Triadic interactions are known to be important in various systems, including: ecosystems [25, 26, 27] where one species can regulate the interaction between two other species; neuronal networks [28], where glial cells regulate synaptic transmission between neurons thereby regulating brain information processing; and gene regulation networks [29, 30], where a modulator can promote or inhibit the interaction between a transcription factor and a target gene. There is mounting evidence that triadic interactions can induce (so far largely unexplored) collective phenomena and dynamical states that can reveal important aspects of complex systems [19, 20, 21, 22, 23, 24, 31, 32]. An important advance in this line of research is triadic percolation [19, 20, 21], a theoretical framework that captures the non-trivial dynamics of the giant component. Moreover, recent results demonstrate that triadic interactions can have significant effects on stochastic dynamics [24] and learning  [22, 23]. However, despite the increasing attention that higher-order interactions are receiving, the detection of triadic interactions from network data and node timeseries, is an important scientific challenge that has not been thoroughly explored [33, 34, 29].
@@
In this article, we formulate a dynamical model in which node variables are affected by triadic interactions, and we propose an information theoretic approach, leading to the Triadic Interaction Mining (TRIM) algorithm, for mining triadic interactions. The proposed dynamical model provides evidence of generative mechanisms by which a triadic interaction can induce a significant variability of the mutual information between two nodes at the end-points of an edge. The TRIM algorithm leverages this finding and mines triadic interactions using knowledge of the network structure and the dynamical variables associated to the nodes. The significance of each putative triadic interaction is then assessed using cross-validation with two distinct null models.
@@
In this way, the TRIM algorithm can go beyond monotonicity assumptions regarding the functional form of the regulation of the two linked nodes by the third node (which is at the foundation of previously proposed methods [29]) allowing for broader applications. Significant node triples are also associated with an entropic score function $S$ that grows with the spread of the conditional joint distributions functions of the variables at the two end point of the regulated edge. We test the TRIM algorithm on a benchmark dynamical model, demonstrating its efficiency in detecting true triadic interactions. We also use the TRIM model to mine triadic interactions from gene-expression, to identify ‘trigenic’ processes [35]. We demonstrate that the TRIM algorithm is able to detect known interactions as well as propose a set of new candidate interactions that can then be validated experimentally.
@@
Figure 1: (Panel a) A triadic interaction occurs when a node $Z$, called a regulator node, regulates (either positively or negatively) the interaction between two other nodes $X$ and $Y$. The regulated edge can be conceptualised as a factor node (shown here as a cyan diamond). (Panel b) A network with triadic interactions can be seen as a network of networks formed by a simple structural network and by a bipartite regulatory network between regulator nodes and regulated edges (factor nodes).
@@ II Triadic interactions
III The continuous triadic interaction model
@@
$L_{ij}$ =$=\begin{cases}-a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \sum_{j}a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j,\end{cases}$ (1)
where ${\bf A}=(a_{ij})$ is the adjacency matrix of the network, and $J>0$ is a coupling constant. In the absence of triadic interactions, we assume that the dynamics of the network is associated with a Gaussian process implemented as the Langevin equation
$d\mbox{\boldmath$X$}=-\frac{\delta\mathcal{H}}{\delta\mbox{\boldmath$X$}}+\Gamma\bm{\eta}(t),$ (2)
with the Hamiltonian
$\mathcal{H}=\frac{1}{2}{\mbox{\boldmath$X$}}^{\top}({\bf L}+\alpha{\bf I})\mbox{\boldmath$X$},$ (3)
where $\Gamma>0$, $\alpha>0$, and ${\bm{\eta}}(t)$ indicates uncorrelated Gaussian noise with
$\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}),$ (4)
for all $t$ and $t^{\prime}$. The resulting Langevin dynamics are given by
$d\mbox{\boldmath$X$}=-(\mathbf{L}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}).$ (5)
@@
$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij}.$ (6)
In other words, from the correlation matrix it is possible to infer the Laplacian, and hence the connectivity of the network.
We now introduce triadic interactions in the model. As explained earlier, a triadic interaction occurs when one or more nodes modulate the interaction between another two nodes. To incorporate triadic interactions into the network dynamics, we modify the definition of the Laplacian operator present in the Langevin equation. Namely, we consider the Langevin dynamics
$d\mbox{\boldmath$X$}=-(\mathbf{L}^{(\text{T})}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}),$ (7)
$L^{(\text{T})}_{ij}$ =$=\begin{cases}-a_{ij}J_{ij}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \displaystyle\sum_{k=1}^{N}a_{ik}J_{ik}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j.\end{cases}$ (8)
Moreover, we assume that the coupling constants $J_{ij}(\mbox{\boldmath$X$})$ are determined by a perceptron-like model that considers all the regulatory nodes of the link $\ell=[i,j]$ and the sign of the regulatory interactions. Specifically, if $\sum_{k=1}^{N}K_{\ell k}X_{k}\geq\hat{T}$ then we set $J_{ij}=w^{+}$; if instead $\sum_{k=1}^{N}K_{\ell k}X_{k}<\hat{T}$ then we set $J_{ij}=w^{-}$, with $w_{+},w_{-}\in{}_{+}$ and $w_{+}>w_{-}$. Thus, we set
$J_{ij}(\mbox{\boldmath$X$})$ $=$ $w_{-}+(w_{+}-w_{-})\theta\left(\displaystyle\sum_{k=1}^{N}K_{\ell k}X_{k}-\hat{T}\right),$ (9)
[…]
Our model of triadic interactions, therefore, has the crucial property that triadic interactions are higher-order interactions that cannot be reduced to pairs of pairwise interactions. An important problem that then arises is whether such interactions can be mined from observational data. To address this issue, we will develop a new algorithm – that we call the TRIM algorithm – to identify triadic interactions from data, and we will test its performance on the data generated from the model described above.
Figure 2: The TRIM algorithm identifies triples of nodes $X$, $Y$, and $Z$ involved in a putative triadic interaction, starting from the knowledge of the structural network and the dynamical variables associated with its nodes. For each putative triple of nodes involved in a triadic interaction (panel (a)) which belong to a network whose structure and dynamics is known (panel (b)), we study the functional behavior of the conditional mutual information $MIz$ (panel (c)), and assess the significance of the observed modulations of $MIz$ with respect to a null model (panel (d)). Given a predetermined confidence level, we can use these statistics to identify significant triadic interactions (panel (e)). This procedure can be extended to different triples of the network, thereby identifying the triadic interactions present in it (panel (f)).
@@ IV.1 The TRIM algorithm
Given a structural edge between nodes $X$ and $Y$, our goal is to determine a confidence level for the existence of a triadic interaction involving an edge between node $X$ and node $Y$ with respect to a potential regulator node $Z$. Specifically, we aim to determine whether the node $Z$ regulates the edge between node $X$ and node $Y$, given the dynamical variables $X$, $Y$ and $Z$ associated with these nodes. To do so, given a time series associated with node $Z$, we first sort the $Z$-values, and define $P$ bins in terms of the quantiles of $z$, chosen in such a way that each bin $m_{z}$ comprises the same number of data points (ranging in our analyses from $30$ to $100$). We indicate with $z_{m}$ the quantile of $Z$ corresponding to the percentile $m/P$. Therefore, each bin $m_{z}$ indicates data in which $Z$ ranges in the interval $[z_{m},z_{m+1})$. We indicate with $\mu(x|z_{m}),\mu(y|z_{m})$ and $\mu(x,y|z_{m})$ the probability density of the variables $X$, $Y$ and the joint probability density of the variables $X$ and $Y$ in each $m_{z}$ bin.
@@
[…]
In order to estimate this quantity, we rely on non-parametric methods based on entropy estimation from $k$-nearest neighbors [41, 42, 43]. For each triple of nodes, we visualize the mutual information $MIz$ computed as a function of the $m/P$-th quantiles $z_{m}$ and fit this function with a decision tree comprising $r$ splits.
In the absence of triadic interactions, we expect $MIz$ to be approximately constant as a function of the $m/P$-th quantiles $z_{m}$, while, in the presence of triadic interactions, we expect this quantity to vary significantly as a function of $z_{m}$. The discretized conditional mutual information CMI between $X$ and $Y$ conditioned on $Z$ can be written as
@@
[…]
where $\mu(x),\mu(y)$ are the probability density functions for $X$ and $Y$ and $\mu(x,y)$ is the joint probability density functions of the variables $X$ and $Y$. In particular, true interactions between $X$ and $Y$ are typically associated with both high mutual information MI and high conditional mutual information CMI.
The conditional mutual information CMI, however, is not sensitive to variations in $MIz$ and does not therefore provide the information needed to detect triadic interactions. In order to overcome this limitation, we define the following three quantities that measure how much the mutual information between $X$ and $Y$ conditioned on $Z\in[z_{m},z_{m+1})$ changes as $z_{m}$ varies. Specifically we consider:
@@
    the difference $T$ between the maximum and minimum value of $MIz$, given by
@@
The quantities $\Sigma$, $T$, collectively measure the strength of the triadic interaction under question and can thus be used to mine triadic interactions in synthetic as well as in real data. In order to assess the significance of the putative triadic interaction, we compare the observed values of these variables to the results obtained with given null models. In order to determine if the observed values are significant with respect to a given null model, we compute the scores $\Theta_{\Sigma}$, $\Theta_{T}$, given by
@@
Note that if we consider $\mathcal{N}$ realizations of the null model, we cannot estimate probabilities smaller than $1/\mathcal{N}$. Therefore, if in our null model we observe no value of $\Sigma^{\textup{ran}}$ larger than the true data $\Sigma$, we set the conservative estimate $p_{\Sigma}=1/\mathcal{P}$. A similar procedure is applied also to $p_{T}$.
@@
In this work we have considered two types of null models. The first is the randomization null model obtained by shuffling the $Z$ values, to give $\mathcal{N}$ randomization of the data, i.e. we use surrogate data for testing [44, 45]. The second is the maximum likelihood Gaussian null model between the three nodes involved in the triple $X$,$Y$,$Z$. We note that the adoption of these two null models allows us to also potentially capture non-monotonic relationships between $MIz$ and $z$, thereby going beyond underlying monotonic assumptions made in [46]. The first null model disrupts the temporal correlations between the timeseries of the node $Z$ and the timeseries the two nodes $X$ and $Y$ at the endpoints of the considered edges. Therefore this null model is robust with respect to the presence of possible outliers in the dataset. However this first null model may overlook confounding network effects that affect correlations between the dynamical variables. The second null model more efficiently captures correlations between the dynamical state of the three considered nodes due to network effects but is more sensitive to the presence of outliers in the data. To increase confidence, we therefore combine the insights coming from both null models (see SI for details).
@@
For each triple, the function $MIz(m)$ is fitted with a decision tree with two splits. In this way, three different intervals of values of $Z$ are identified, each corresponding to a distinct functional behavior of the correlation functions between the variables $X$ and $Y$. While the method would allow for more than two splits of the decision three in general, here we have chosen three splits as this is the minimum number of splits that will allow us to capture non-trivial functional behaviour of $MIz$, like non-monotonicity. In practice this will be the best choice in presence of limited data and a small number of bins as in the Gene Expression data we will analyse in the following.
@@
We further characterize significant triples by examining their entropic score function $S$, which is used to characterize their corresponding functional behaviour, as described in the SI.
@@
As we will discuss below, the algorithm performs well on our paradigmatic model. Moreover, while the results obtained with the randomized surrogate data might neglect potentially important network effects, this shortcoming of this approach is mitigated by the cross-validation with the Gaussian null model.
@@
Figure 3: Representative results for triples of nodes involved in triadic interactions in the continuous model with triadic interactions. The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ are shown in panel (a). Panel (b) shows the behavior of $MIz$ as a function of the values of $z_{m}$, which clearly departs from the constant behavior expected in absence of triadic interactions. Panel (c) presents the decision tree for fitting the $MIz$ functional behavior and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditioned on $Z$ are observed. For more details about the simulation results, see the Supplementary Information (SI).
@@
Figure 4: We consider a network with $N=10$, nodes $L=12$ edges, and $\hat{L}=5$ regulatory interactions (panel (a)). The time series obtained by integrating the stochastic dynamics of the proposed dynamical model for triadic interactions (Eq.(7)) are analyzed with the TRIM algorithm. Panel (b) displays the Receiver Operating Characteristic curve (ROC curve) obtained by running TRIM with $P=400$ bins and $\mathcal{N}=10^{3}$ realizations of the null model on these synthetic time series, using $\Theta_{\Sigma}$ to score for different parameters values indicated in the legend. The time series is simulated up to a maximum time $t_{max}=4000$ with a $dt=10^{-2}$. For the analysis, we consider $40000$ time steps (see the SI for details). The parameters of the model are: $\hat{T}=10^{-3}$, $w^{+}=8$, $w^{-}=0.5$, and $\alpha$ and $\Gamma$ as indicated in the figure legend.
IV.2 Validation of the TRIM algorithm on the triadic interaction model

### v4

@@ Abstract
Complex systems often involve higher-order interactions which require us to go beyond their description in terms of pairwise networks. Triadic interactions are a fundamental type of higher-order interaction that occurs when one node regulates the interaction between two other nodes. Triadic interactions are found in a large variety of biological systems, from neuron-glia interactions to gene-regulation and ecosystems. However, triadic interactions have so far been mostly neglected. In this article, we propose the Triadic Perceptron Model (TPM) that demonstrates that triadic interactions can modulate the mutual information between the dynamical state of two linked nodes. Leveraging this result, we formulate the Triadic Interaction Mining (TRIM) algorithm to extract triadic interactions from node metadata, and we apply this framework to gene expression data, finding new candidates for triadic interactions relevant for Acute Myeloid Leukemia. Our work reveals important aspects of higher-order triadic interactions that are often ignored, yet can transform our understanding of complex systems and be applied to a large variety of systems ranging from biology to climate.
@@ I Introduction
Higher-order networks [1, 2, 3, 4, 5] are key to capturing many-body interactions present in complex systems. Inferring higher-order interactions [6, 7, 8, 9, 10, 11] from real, pairwise network datasets is recognised as one of the central challenges in the study of higher-order networks [2, 12], with wide applicability across different scientific domains, from biology and brain research [13, 14, 15] to finance [16, 17]. Mining higher-order interactions from the exclusive knowledge of the pairwise networks typically involves generative models and Bayesian approaches based on network structural properties [6, 18, 7, 9, 11]. Note, however, that when the inference is performed on the basis of the knowledge of the nodes’ dynamical states [10, 8], inferring higher-order interactions also requires dynamical considerations.
@@
Triadic interactions [19] are a fundamental type of signed higher-order interaction that are gaining increasing attention from the statistical mechanics community [19, 20, 21, 22, 23, 24], since they are not reducible to hyperedges or simplices. A triadic interaction occurs when one or more nodes regulate the interaction between two other nodes. The regulator nodes may either enhance or inhibit the interaction between the other two nodes. Triadic interactions are known to be important in various systems, including: ecosystems [25, 26, 27] where one species can regulate the interaction between two other species; neuronal networks [28], where glial cells regulate synaptic transmission between neurons thereby controlling brain information processing; and gene regulatory networks [29, 30], where a modulator can promote or inhibit the interaction between a transcription factor and its target gene. There is mounting evidence that triadic interactions can induce collective phenomena and/or modulate dynamical states that reveal important aspects of complex system behavior [19, 20, 21, 22, 23, 24, 31, 32]. An important advance in this line of research is triadic percolation [19, 20, 21], a theoretical framework that captures the non-trivial dynamics of the giant component. Moreover, recent results demonstrate that triadic interactions can have significant effects on stochastic dynamics [24] and learning [22, 23]. However, despite the increasing attention that higher-order interactions are receiving, the detection of triadic interactions from network data and node time series, is an important scientific challenge that has not been thoroughly explored [33, 34, 29].
@@
In this article, we formulate the Triadic Perceptron Model (TPM) in which continuous node variables are affected by triadic interactions. Based on the insights gained by investigating this model, we propose an information theoretic approach, leading to the Triadic Interaction Mining (TRIM) algorithm, for mining triadic interactions. The TPM provides evidence of the mechanisms by which a triadic interaction can induce a significant variability of the mutual information between two nodes at the end-points of an edge. The TRIM algorithm leverages this finding and mines triadic interactions using knowledge of the network structure and the dynamical variables associated with the nodes. The significance of each putative triadic interaction is then validated by comparison with two distinct null models.
@@
In this way, the TRIM algorithm can go beyond monotonicity assumptions regarding the functional form of the regulation of the two linked nodes by the third node (which is at the foundation of previously proposed methods [29]) allowing for broader applications. Significant node triples are also associated with an normalized entropic score function $S\in[0,1]$ that quantifies the spread of the conditional joint distribution functions of the variables at the ends of the regulated edge. We test the TRIM algorithm on the benchmark TPM, demonstrating its efficiency in detecting true triadic interactions. We also use the TRIM model to mine triadic interactions from gene-expression, to identify ‘trigenic’ processes [35]. We demonstrate that the TRIM algorithm is able to detect known interactions as well as propose a set of new candidate interactions that can then be validated experimentally.
@@
Figure 1: (Panel a) A triadic interaction occurs when a node $Z$, called a regulator node, regulates (either positively or negatively) the interaction between two other nodes $X$ and $Y$. The regulated edge can be conceptualized as a factor node (shown here as a cyan diamond). (Panel b) A network with triadic interactions can be seen as a network of networks formed by a simple structural network and by a bipartite regulatory network between regulator nodes and regulated edges (factor nodes).
@@ II Triadic interactions
III The Triadic Perceptron Model (TPM)
Here, we formulate a model for node dynamics in a network with triadic interactions that we call Triadic Perceptron Model (TPM). The TPM acts as a benchmark to validate the TRIM algorithm proposed here. We assume that each node $i$ of the network is associated with a dynamical variable $X_{i}\in\mathbb{R}$, and that the dynamical state of the entire network is encoded in the state vector $\mbox{$X$}=(X_{1},X_{2}\ldots,X_{N})^{T}$. The topology of the structural network is encoded in the graph Laplacian matrix $\bf L$} with elements
$L_{ij}$ $=\begin{cases}-a_{ij}J&\text{if}\penalty 10000\ \penalty 10000\ i\neq j,\\[2.5pt] \sum_{k}a_{ik}J&\text{if}\penalty 10000\ \penalty 10000\ i=j,\end{cases}$ (1)
where $\bf a$} is the adjacency matrix of the network of elements $a_{ij}$, and $J>0$ is a coupling constant. In the absence of triadic interactions, we assume that the dynamics of the network is associated with a Gaussian process implemented as the Langevin equation
$\frac{d\mbox{$X$}}{dt}=-\frac{\delta\mathcal{H}}{\delta\mbox{$X$}}+\Gamma\bm{\eta}(t),$ (2)
with the Hamiltonian
$\mathcal{H}=\frac{1}{2}{\mbox{$X$}}^{\top}({\bf L}+\alpha{\bf I})\mbox{$X$},$ (3)
where $\Gamma>0$, $\alpha>0$, and where ${\bm{\eta}}(t)$ indicates uncorrelated Gaussian noise with
$\left\langle{\eta_{i}(t)}\right\rangle=0,\quad\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}),$ (4)
for all $t$ and $t^{\prime}$. The resulting Langevin dynamics are given by
$\frac{d\mbox{$X$}}{dt}=-(\mathbf{L}+\alpha\mathbf{I})\mbox{$X$}+\Gamma\bm{\eta}({t}).$ (5)
We remark that the Hamiltonian $\mathcal{H}$ has a minimum for $\mbox{$X$}={\bf 0}$, and its depth increases as the value of $\alpha$ increases. In a deterministic version of the model ($\Gamma=0$), the effect of the structural interactions will not be revealed at stationarity. However, the Langevin dynamics with $\Gamma>0$ encode the topology on the network. Indeed, at equilibrium the correlation matrix $C_{ij}=\mathbb{E}((X_{i}-\mathbb{E}(X_{i}))(X_{j}-\mathbb{E}(X_{j})))$ is given by
$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij},$ (6)
see Supplementary Information (SI) for details. In other words, from the correlation matrix it is possible to infer the Laplacian, and hence the connectivity of the network.
We now introduce triadic interactions in the TPM. As explained earlier, a triadic interaction occurs when one or more nodes modulate the interaction between another two nodes. To incorporate triadic interactions into the network dynamics, we modify the definition of the Laplacian operator present in the Langevin equation. Namely, we consider the Langevin dynamics
$\frac{d\mbox{$X$}}{dt}=-(\mathbf{L}^{(\text{T})}+\alpha\mathbf{I})\mbox{$X$}+\Gamma\bm{\eta}({t}),$ (7)
$L^{(\text{T})}_{ij}$ $=\begin{cases}-a_{ij}J_{ij}(\mbox{$X$})&\text{if}\penalty 10000\ \penalty 10000\ i\neq j,\\[2.5pt] \displaystyle\sum_{k=1}^{N}a_{ik}J_{ik}(\mbox{$X$})&\text{if}\penalty 10000\ \penalty 10000\ i=j.\end{cases}$ (8)
Moreover, we assume that the coupling constants $J_{ij}(\mbox{$X$})$ are determined by a perceptron-like model that considers all the regulatory nodes of the link $\ell=[i,j]$ and the sign of the regulatory interactions. Specifically, if $\sum_{k=1}^{N}K_{\ell k}X_{k}\geq\hat{T}$ then we set $J_{ij}=w^{+}$; if instead $\sum_{k=1}^{N}K_{\ell k}X_{k}<\hat{T}$ then we set $J_{ij}=w^{-}$, with $w_{+},w_{-}\in\real_{+}$ and $w_{+}>w_{-}$. Thus,
@@ III The Triadic Perceptron Model (TPM)
The TPM for continuous node dynamics in presence of triadic interactions is very general and comprehensively expresses the modulation of structural interactions by other nodes in the network. Therefore, the dynamics of TPM cannot be reduced to dynamics exclusively determined by pairwise interactions. An important problem that then arises is whether such interactions can be mined from observational data. To address this issue, we will develop a new algorithm – that we call the TRIM algorithm – to identify triadic interactions from data, and we will test its performance on the data generated from the TPM model described above.
Figure 2: The TRIM algorithm identifies triples of nodes $X$, $Y$, and $Z$ involved in a putative triadic interaction, starting from the knowledge of the structural network and the dynamical variables associated with its nodes. For each putative triple of nodes involved in a triadic interaction (panel (a)) which belong to a network whose structure and dynamics is known (panel (b)), we study the functional behavior of the conditional mutual information MIzMIz (panel (c)), and assess the significance of the observed modulations of MIzMIz with respect to a null model (panel (d)). Given a predetermined confidence level, we can use these statistics to identify significant triadic interactions (panel (e)). This procedure can be extended to different triples of the network, thereby identifying the triadic interactions present in it (panel (f)).
@@ IV.1 The TRIM algorithm
Given a structural edge between nodes $X$ and $Y$, our goal is to determine a confidence level for the existence of a triadic interaction involving an edge between node $X$ and node $Y$ with respect to a potential regulator node $Z$. Specifically, we aim to determine whether the node $Z$ regulates the edge between node $X$ and node $Y$, given the dynamical variables $X$, $Y$ and $Z$ associated with these nodes. To do so, given a time series associated with node $Z$, we first sort the $Z$-values, and define $P$ bins in terms of the quantiles of $z$, chosen in such a way that each bin $m_{z}$ comprises the same number of data points (ranging in our analyses from $30$ to $100$). We indicate with $z_{m}$ the quantile of $Z$ corresponding to the percentile $m/P$. Therefore, each bin $m_{z}$ indicates data in which $Z$ ranges in the interval $z_{m},z_{m+1})$. We indicate with $\mu(x|z_{m}),\mu(y|z_{m})$ and $\mu(x,y|z_{m})$ the probability density of the variables $X$, $Y$ and the joint probability density of the variables $X$ and $Y$ in each $m_{z}$ bin.
@@
$MIz(m)=\int dx\int dy\mu(x,y|z_{m})\log\left(\frac{\mu(x,y|z_{m})}{\mu(x|z_{m})\mu(y|z_{m})}\right).$
In order to estimate this quantity, we rely on non-parametric methods based on entropy estimation from $k$-nearest neighbours [41, 42, 43] (see SI for details). For each triple of nodes, we visualize the mutual information $MIz$ computed as a function of the $m/P$-th quantiles $z_{m}$ and fit this function with a decision tree comprising $r$ splits.
In the absence of triadic interactions, we expect $MIz$ to be approximately constant as a function of the $m/P$-th quantiles $z_{m}$, while, in the presence of triadic interactions, we expect this quantity to vary significantly as a function of $z_{m}$. The discretised conditional mutual information CMI between $X$ and $Y$ conditioned on $Z$ can be written as
@@
$MI_{X,Y}=\int dx\int dy\mu(x,y)\log\left(\frac{\mu(x,y)}{\mu(x)\mu(y)}\right),$
where $\mu(x),\mu(y)$ are the probability density functions for $X$ and $Y$ and $\mu(x,y)$ is the joint probability density functions of the variables $X$ and $Y$. The conditional mutual information, however, is not sensitive to variations in $MIz$ and does not therefore provide the information needed to detect triadic interactions. In order to overcome this limitation, we define the following two quantities that measure how much the mutual information between $X$ and $Y$ conditioned on $Z\in[z_{m},z_{m+1})$ changes as $z_{m}$ varies. Specifically we consider:
@@
    the difference $T$ between the maximum and average value of $MIz$, given by
@@
The quantities $\Sigma$ and $T$, collectively measure the strength of the triadic interaction under question and can thus be used to mine triadic interactions in synthetic as well as in real data. In order to assess the significance of the putative triadic interaction, we compare the observed values of these variables to the results obtained with given null models. In order to determine if the observed values are significant with respect to a given null model, we compute the scores $\Theta_{\Sigma}$, $\Theta_{T}$, given by
@@
Note that if we consider $\mathcal{N}$ realizations of the null model, we cannot estimate probabilities smaller than $1/\mathcal{N}$. Therefore, if in our null model we observe no value of $\Sigma^{\textup{ran}}$ larger than the true data $\Sigma$, we set the conservative estimate $p_{\Sigma}=1/\mathcal{N}$. A similar procedure is applied also to $p_{T}$.
@@
To assess this significance we consider two types of null models. The first is the randomization null model obtained by shuffling the $Z$ values, to give $\mathcal{N}$ randomization of the data, i.e. we use surrogate data for testing [44, 45]. The second is the maximum likelihood Gaussian null model between the three nodes involved in the triple $X$,$Y$,$Z$. Specifically, the Gaussian null model uses the mean and covariance of the timeseries of $X,Y$ and $Z$ to define a multivariate normal distribution from which samples are randomly drawn, thereby providing surrogate timeseries values for the considered triple. We note that the use of these two null models allows us to identify also non-monotonic relationships between $MIz$ and $z$, thereby going beyond underlying monotonic assumptions made elsewhere [46]. The first null model disrupts the temporal correlations between the timeseries of the node $Z$ and the timeseries the two nodes $X$ and $Y$ at the endpoints of the considered edges. Therefore this null model is robust with respect to the presence of possible outliers in the dataset. However, this first null model may overlook confounding network effects that affect correlations between the dynamical variables. The second null model more efficiently captures correlations between the dynamical state of the three considered nodes due to network effects but is more sensitive to the presence of outliers in the data. To increase confidence, we therefore combined the insights coming from both these null models (see SI for details).
@@
For each triple, the function $MIz(m)$ is fitted with a decision tree with two splits. In this way, three different intervals of values of $Z$ are identified, each corresponding to a distinct functional behavior of the correlation functions between the variables $X$ and $Y$. While our method in principle allows for more than two splits of the decision tree, for illustrative purposes we have chosen two splits since this is the minimum number of splits needed to capture non-trivial functional behavior in $MIz$, such as non-monotonicity. In practice this choice of two splits will also be the best choice when data is limited, such as the gene expression data we will analyze in the following section.
@@
We also further characterize significant triples by examining their normalized entropic score function $S\in[0,1]$, which is used to characterize their corresponding functional behavior. Specifically, th entropic score $S$ classifies the diversity of each of the joint distribution functions of $X$ and $Y$ conditioned on $Z$ for each interval obtained through the decision tree (see SI for details).
@@
As we will discuss below, the algorithm performs well on data obtained from the TPM. In this case, we also observe that true triadic triples are characterized by a high entropic score $S$. On real data, the results obtained with the TRIM algorithm using randomized surrogate data might neglect potentially important network effects, this shortcoming is mitigated by performing an additional validation using the Gaussian null model and the entropic score $S$ (see SI for the full pipeline of TRIM).
Figure 3: We consider a network with $N=10$ nodes, $L=12$ edges, and $\hat{L}=5$ triadic interactions (panel (a)). Panels (b) and (c) display the effect of triadic interactions on the Mutual Information profile MIzMIz. Panel (b) shows MIzMIz for the triple $[4,9,5]$ involved in a positive triadic interaction. Panel (c) shows MIzMIz for the triple $[1,2,6]$ that is not involved in a triadic interaction. In all panels simulations were run to $t_{\textup{max}}=4,000$ with a timestep of $dt=10^{-2}$. For the analysis we consider $40,000$ time steps. The parameters of the model are: $\alpha=0.05,\hat{T}=10^{-3},\Gamma=10^{-2}$, $w^{+}=8,w^{-}=0.5$, number of bins $P=400$.
@@
Figure 4: Representative results for triples of nodes involved in triadic interactions in the continuous model with triadic interactions. Results for the triple $[4,9,5]$ of the network in Figure 3 of the main text, which is triadic, are shown. The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ are shown in panel (a). Panel (b) shows the behavior of MIzMIz as a function of the values of $z_{m}$, which clearly departs from the constant behavior expected in absence of triadic interactions. Panel (c) presents the decision tree for fitting the $MIz$ functional behavior and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditioned on $Z$ are observed. The parameters used are the same as in Figure 3.
@@
Figure 5: We consider the network in Figure 3(a). The time series obtained by integrating the stochastic dynamics of the proposed dynamical model for triadic interactions (Eq.(7)) are analyzed with the TRIM algorithm. Panel (a) displays the Receiver Operating Characteristic curve (ROC curve) obtained by running TRIM with $P=400$ bins and $\mathcal{N}=10^{3}$ realizations of the null model on these synthetic time series, using $\Theta_{\Sigma}$ to score for different parameters values indicated in the legend. Panel (b) displays the corresponding Precision-Recall curve (PR curve) obtained by running TRIM with the same parameters. The timeseries are simulated up to a maximum time $t_{max}=4000$ with a $dt=10^{-2}$. For the analysis, we consider $40,000$ time steps (see the SI for details). The parameter of the model are: $\hat{T}=10^{-3}$, $w^{+}=8$, $w^{-}=0.5$, and $\alpha$ and $\Gamma$ as indicated in the figure legend.
IV.2 Validation of the TRIM algorithm on the triadic perceptron model